What is 3L=ml
what is 70cm= mm
what is 3m=dm

Answer by RAY100(1637)

3L *(1000ml/1L) = 3000 ml
.
70 cm * (1m/100cm) * (1000mm/1m) = 700 mm
.
3m * (10 dm / 1 m ) = 30 dm
.